Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probablility PS

Expert replies
by theachiever » Tue Jul 16, 2013 3:19 am
Gender Under 2.0 2.0-3.0 Over 3.0

Male 0.05 0.25 0.10

Female 0.10 0.30 0.20

•If the student selected is female, what is the probability that her GPA is between 2.0 and 3.0?
•If the GPA of the student selected is over 3.0, what is the probability that the student is male?
•What is the probability that the student selected is female or has a GPA under 2.0 or both?
• Is GPA independent of gender?
Resilience is a trait that manifests in an individual when he/she faces consistent failures motivating them to succeed in the end.
Join the discussion
Source: — Problem Solving |

by faraz_jeddah » Tue Jul 16, 2013 4:28 am
I believe that the question is in a table format (3x3)


if the student selected is female, what is the probability that her GPA is between 2.0 and 3.0?
Answer - 0.3

If the GPA of the student selected is over 3.0, what is the probability that the student is male?
Answer - 0.1 / (0.1 + 0.2) = 1/3 = 0.33

What is the probability that the student selected is female or has a GPA under 2.0 or both?

Answer = 0.6 + 0.15 + 0.1 = 0.85

Is GPA independent of gender?
Yes ?
Join the discussion

by jitsy » Tue Jul 16, 2013 7:37 am
This involves conditional probability. The probability of A conditioned on B, denoted P(A|B), is equal to P(AB)/P(B). The division provides that the probabilities of all outcomes within B will sum to 1. Conditioning restricts the sample space to those outcomes which are in the set being conditioned on (in this case B).

Answering question 1

1) If the student selected is female, what is the probability that her GPA is between 2.0 and 3.0?

Here
P(A)= 0.3/0.6 = 0.5
P(B)= 0.6/1.0 = 0.6
P(A|B)= P(AB)/P(B)= P(A)*P(B)/P(B) = 0.5*0.6/0.6= 0.5
Join the discussion

by theachiever » Tue Jul 16, 2013 9:42 am
Thanks for the solutions guys.I appreciate the help

Experts can you please give your inputs about the problem if you can guide me about it,it would be helpful.
Resilience is a trait that manifests in an individual when he/she faces consistent failures motivating them to succeed in the end.
Join the discussion

by Matt@VeritasPrep » Tue Jul 16, 2013 10:23 am
Let's do it!

The first question is asking what percentage of the females have a GPA of 3.0, so our probability is:

# of 3.0 females
----------------
Total # of females

= .3/.6 = .5 = 50%

The second question is similar

# of males over 3.0
-------------------
Total # of people over 3.0

= .1/.3 = 1/3 = 33.3..%

The third one is tricky: for the odds of AT LEAST ONE of two groups, we do (Odds A) + (Odds B) - (Odds of A + B).

So ...

Odds Female = .6/1 = 60%
Odds Under 2.0 = .15/1 = 15%
Odds Both = .1/1 = 10%

So 60% + 15% - 10% = 65%

The other way to find the odds of AT LEAST ONE of two things is to do 1 - (Odds of neither). So the odds that someone is NOT a female or a person with a GPA under 2.0 is .35 (Guys with 2.0-3.0 or Guys with 3.0+). So 1 - .35 = .65, or 65%.
Join the discussion

by jitsy » Thu Jul 18, 2013 9:38 am
Matt,

Could you please help me with the last part of the (Yes/No) question and also explaining the reasoning. Thank you. I think it is the same as the last part in the question below. Just need some guidance in answering questions like these.

On 43percent of days, the market actually goes up. On 12 percent of days, Peter predicts that the market will go up and it does. If we pick a day at random, what is the probability that:
•Peter will predict that the market will go up, or it will actually go up, or both?
•The market will actually go up given that Peter has predicted it will?
•Does the Stock market moves up independently of Peter’s Prediction?

Matt@VeritasPrep wrote:Let's do it!

The first question is asking what peurcentage of the females have a GPA of 3.0, so our probability is:

# of 3.0 females
----------------
Total # of females

= .3/.6 = .5 = 50%

The second question is similar

# of males over 3.0
-------------------
Total # of people over 3.0

= .1/.3 = 1/3 = 33.3..%

The third one is tricky: for the odds of AT LEAST ONE of two groups, we do (Odds A) + (Odds B) - (Odds of A + B).

So ...

Odds Female = .6/1 = 60%
Odds Under 2.0 = .15/1 = 15%
Odds Both = .1/1 = 10%

So 60% + 15% - 10% = 65%

The other way to find the odds of AT LEAST ONE of two things is to do 1 - (Odds of neither). So the odds that someone is NOT a female or a person with a GPA under 2.0 is .35 (Guys with 2.0-3.0 or Guys with 3.0+). So 1 - .35 = .65, or 65%.
If you appreciated my post or even just my time trying to help, please don't forget to click on 'Thanks' and say "Piece o' Cake mate" at the same time.
Join the discussion